OpenStudy (anonymous):

y is directly proportional to x^2. It is known that y=10 for a particular value of x. Find the value of y when the value of x is halved

OpenStudy (anonymous):

y = k*x*x agree? so 10 = k*x*x k=10/x/x so y= 10/x/x * (x/2)*(x/2)

OpenStudy (anonymous):

10/4